# An Efficient Instance Segmentation Framework Based on Oriented Bounding Boxes

### v3

@@ III-C Knowledge Distillation
BSMs exhibit powerful performance, however they usually have high computational complexity. Some methods, such as MobileSAM [19], focus on distilling image encoders. However, the performance gap between the distilled model and the teacher model is still large. To tackle this, CFNet uses the distilled image encoders, and further distills the prompt encoders and mask decoders of BSMs. Considering the powerful data-fitting capability of the teacher model, the student model only imitates the teacher model and does not learn from GT labels. Furthermore, to further enhance the performance and generalization capability of the student model, Gaussian smoothing is applied to teacher model outputs. The process of knowledge distillation for the proposed OBB prompt encoder and mask decoder is shown in Fig. 4.
We define a one-dimensional Gaussian kernel $\mathbf{G}^{1}_{k}(x_{i};\sigma)$ with a length of $k$, which is given as
@@ IV-A Dataset
To verify the effectiveness of the proposed methods, we conduct completely occluded object instance segmentation and dense object instance segmentation experiments on self-collected industrial and public datasets, respectively.
Figure 5: Self-designed robotic system for completely occluded object instance segmentation in the industrial robot assembly environment of the large commercial aircraft C919.
IV-A1 Industrial Dataset for Completely Occluded Objects
The self-designed industrial robotic vision system for large commercial aircraft assembly is shown in Fig. 5. It mainly consists of a UR10 robot, a vision sensor and a computer. Specifically, the UR10 robot is used to carry the vision sensor for operations. The computer is responsible for receiving and processing the image data collected by the vision sensor, and it runs the algorithm for completely occluded object instance segmentation. Due to the limited working space, reference holes are usually observed from the side.
All the images in the dataset are collected by this system, including three types of occluded reference holes, i.e., reference holes with bolts, nuts and untreated bolts (see the left image of Fig. 1). To reflect complex and changeable environments in the assembly process, each type of reference hole is photographed with the camera at various perspectives, distances and lighting conditions. We obtain a total of 800 images, with approximately the same number of each type. These images are divided into a training set, a validation set, and a testing set in a ratio of $7\colon 1\colon 2$ in a uniform sampling manner. All images are cropped to $1024\times 1024$. Then, these images are first flipped horizontally and vertically to quadruple the number. Experimental GT labels are obtained from industrial standard workpieces. Furthermore, the Mosaic method that randomly mixes four images is used for data augmentation.
IV-A2 Public Dataset for Dense Objects
iSAID [8] is a large-scale and densely annotated instance segmentation dataset in aerial images. It has 655,451 object instances across 15 categories. The training set, validation set and testing set contain 1411, 458 and 937 images, respectively. All images are cropped into $1024\times 1024$ patches with a stride of 500. Since iSAID dataset does not provide the GT parameters of OBBs, the smallest OBBs of instance masks are used as regression targets. We only choose densely distributed large vehicles, small vehicles and ships as targets for dense object instance segmentation.
Figure 6: Some visualization results on the completely occluded object instance segmentation dataset. Bounding boxes of different colors and segmentation masks of different colors represent different object classes and different instances, respectively.
@@ IV-B Implementation Details
The training settings for Oriented R-CNN [14] are the same as the original method. In the OBB prompt encoder, the length of the sampled Fourier basis frequencies is 128. For knowledge distillation, the teacher model (called CFNet*) is based on SAM with the powerful ViT (vision transformer)-H image encoder and the proposed OBB prompt encoder. The student model (i.e., CFNet) uses a lightweight pretrained ViT-Tiny image encoder and the proposed OBB prompt encoder. Their mask decoders and OBB detection modules are the same. During training, the weights of their image encoders are frozen, and the other parts are fine-tuned. The teacher model has the same training loss as SAM and has been trained before distillation. The trade-off factor $\lambda$ for the distillation loss is 0.1. More implementation details can be found in our open-source code. All experiments are conducted on an RTX 3080Ti GPU. For evaluation, we use the standard COCO metrics: AP (i.e., AP50-95), AP50 and AP75, where “AP” means average precision over IoU (Intersection over Union) threshold.
IV-C Completely Occluded Object Instance Segmentation
In completely occluded object instance segmentation experiments, we compare CFNet, SAM, and amodal methods. SAM uses the ViT-H to encode images and the HBB detection branch of Mask R-CNN to generate HBB prompts (the same below). Amodal methods directly predict occludee instances. CFNet and SAM perform the proposed instance segmentation method on occluders to obtain occludee instances. We also include SAM after fine-tuning on this dataset and the teacher model (CFNet*) used for distillation in comparison. For fair comparison, we replace the backbones of current amodal instance segmentation methods with the ViT-H image encoder used in BSMs (indicated by $\dagger$).
TABLE I: Comparison results on the completely occluded object instance segmentation dataset.
Method AP AP50 AP75 FPS
AmodalMask [15] 19.49 35.18 24.36 16.1
ASN [5] 40.88 61.24 45.93 18.6
ORCNN [6] 37.60 57.78 40.16 22.2
BCNet [4] 45.30 65.13 51.79 7.4
@@ IV-C Completely Occluded Object Instance Segmentation
SAM [13] 26.15 42.74 30.28 2.8
Fine-tuned SAM 68.66 82.23 71.38 2.8
CFNet* 75.13 89.21 77.98 2.7
CFNet 73.46 86.48 76.20 20.4
As shown in Table I, the accuracy of CFNet and CFNet* is superior to other methods, especially for high-precision metrics. Additionally, CFNet exhibits competitive inference speed. Some visualization results are presented in Fig. 6. Compared with amodal instance segmentation methods, CFNet provides an effective solution for completely occluded object instance segmentation. In addition, compared with fine-tuned SAM using HBBs, the accuracy of CFNet and CFNet* using OBBs is significantly improved, which indicates the necessity of using OBBs. Moreover, our proposed knowledge distillation method significantly reduces the computational complexity with a minor loss in accuracy.
Figure 7: Some visualization results on the dense object instance segmentation dataset iSAID. Bounding boxes of different colors and segmentation masks of different colors represent different object classes and different instances, respectively.
IV-D Dense Object Instance Segmentation
For dense object instance segmentation, we compare CFNet, SAM (with HBBs), and instance segmentation methods using OBBs. We also include fine-tuned SAM and the teacher model used for distillation in comparison, and replace the backbones of current instance segmentation methods using OBBs with the ViT-H image encoder used in BSMs.
TABLE II: Comparison results on the dense object instance segmentation dataset. “RB Mask R-CNN” represents Rotated Blend Mask R-CNN [9].
Method AP AP50 AP75 FPS
OIS [10] 27.41 47.85 30.71 18.4
RB Mask R-CNN [9] 29.07 50.96 32.64 16.3
ISOP [12] 34.73 59.16 36.84 19.6
@@ IV-D Dense Object Instance Segmentation
SAM [13] 14.37 41.88 18.52 2.8
Fine-tuned SAM 20.26 45.30 24.59 2.8
CFNet* 45.10 66.45 47.92 2.7
CFNet 43.76 64.29 46.13 20.4
As presented in Table II, CFNet and CFNet* outperform other methods, and CFNet achieves the highest inference speed of 20.4 FPS. In addition, CFNet achieves a remarkable dense object instance segmentation accuracy of 43.76% AP, 64.29% AP50, and 46.13% AP75. Fig. 7 shows some dense object instance segmentation results.
Compared with existing instance segmentation methods using OBBs, CFNet achieves higher accuracy since it reduces the dependence on OBBs and is based on BSMs. Significant accuracy improvements over fine-tuned SAM also indicate the benefit of using OBBs. Moreover, compared with CFNet*, the notable inference speed improvement and minor accuracy loss of CFNet also demonstrate the effectiveness of the proposed knowledge distillation method.
IV-E Ablation Studies
@@
To evaluate the effectiveness of the components of CFNet, a series of ablation experiments are conducted on the iSAID dataset. We use the most commonly used AP50 as the metric.
IV-E1 Oriented Bounding Box Prompt Encoder
@@
The OBB prompt encoder introduces positional encoding and specific embeddings to encode OBBs. In Table III, “OCE”, “TCE” and “BCE” represent the specific orientation correction embeddings, top-left corner embeddings and bottom-right corner embeddings, respectively. Compared with GPE, AGPE encodes location information better by adaptively learning parameters of Gaussian distributions. Specific embeddings further improve instance segmentation performance.
TABLE III: Performance of each component of the OBB prompt encoder.
@@
61.40
62.27
63.56
64.29
TABLE IV: Comparison of different learning targets and label smoothing methods in knowledge distillation.
@@
63.51 63.72 64.01 64.07 64.15 64.19 64.24
IV-E2 Knowledge Distillation
We first compare the performance of CFNet learning from the teacher model and GT labels, and test the impact of different label smoothing methods. “T” represents learning from the teacher model. The smoothing factor in ULS is 0.1. Gaussian smoothing (GS) has the same settings as SVLS, with a Gaussian kernel size of $3\times 3$ and a standard deviation of 1.0. The weight factor of the teacher model and GT labels is 0.1. “GS-” represents distillation only for the mask decoder.
The experimental results in Table IV demonstrate that learning from the teacher model has higher performance than learning from GT labels (learning only from GT labels means no knowledge distillation is performed), which is mainly thanks to the powerful data-fitting capabilities of BSMs. Furthermore, adding distillation for the OBB prompt encoder effectively enhances performance. Moreover, Gaussian smoothing further improves instance segmentation performance and is superior to ULS. In addition, we compare the impact of different parameter settings in Gaussian smoothing on instance segmentation performance in Table V. When $k=5$, $\sigma=0.3$, and $\delta=1.0$, the highest instance segmentation accuracy of 64.29% AP50 is achieved.
TABLE V: Impact of parameter settings in Gaussian smoothing.
@@ IV-E2 Knowledge Distillation
$k=3$ 64.24 64.24 64.20 64.17
$k=5$ 64.22 64.29 64.18 64.21
$k=7$ 64.19 64.25 64.15 64.14
IV-E3 Oriented Bounding Box Detection

### v6

@@ III-C Knowledge Distillation
BSMs exhibit powerful performance, however they usually have high computational complexity. Although some lightweight BSMs, such as EfficientSAM, MobileSAM, and EdgeSAM, perform knowledge distillation on BSMs, the performance gap between the distilled model and the teacher model is still large. To tackle this, OBSeg introduces a Gaussian smoothing method to enhance distillation performance. Considering the powerful data-fitting capability of the teacher model, the student model only mimics the teacher model and does not learn from GT labels. Furthermore, to further enhance the performance and generalization capability of the student model, Gaussian smoothing is applied to teacher model outputs. While current methods focus on reducing the complexity of the image encoder, we focus on further improving the distillation performance of the OBB prompt encoder and mask decoder. We use a pretrained lightweight image encoder and perform distillation on the OBB prompt encoder and mask decoder, as shown in Fig. 4.
We first define a one-dimensional Gaussian kernel $\mathbf{G}^{1}_{k}(x_{i};\sigma)$ with a length of $k$, which is given as
@@
Figure 5: Some visualization results of OBSeg’s predictions on the iSAID dataset. Segmentation masks of different colors represent different instances.
@@
TABLE I: Comparison results on the iSAID validation set. “RB Mask R-CNN” represents Rotated Blend Mask R-CNN (same elsewhere)
Method Backbone AP AP50 AP75 APS APM APL FPS
SOLO [28] Inception-ResNet 24.7 44.9 24.2 7.9 32.7 47.8 13.2
Box2Mask-C [29] ResNet-101-FPN 26.6 50.6 23.8 10.6 33.6 47.4 11.1
Luo et al. [30] ResNet-50-FPN 29.4 54.5 27.8 15.5 37.8 42.0 10.7
OIS [7] ResNet-50-FPN 34.1 57.2 34.2 18.5 42.1 49.9 11.0
PointRend [31] ResNet-50-FPN 35.6 59.0 37.3 20.3 44.8 52.9 15.7
Mask Scoring R-CNN [32] ResNet-50-FPN 35.9 57.7 38.4 20.8 44.3 51.5 9.8
DCTC [33] DLASeg 36.3 58.2 38.0 21.2 45.2 47.3 23.9
SCNet [34] ResNet-50-FPN 37.3 59.5 40.3 23.3 44.8 52.3 7.6
Hybrid Task Cascade [35] ResNet-50-FPN 37.4 60.2 40.1 23.5 44.6 53.5 5.8
Ye et al. [36] ResNet-50-FPN 37.6 60.8 41.4 22.4 47.7 53.4 5.3
RB Mask R-CNN [6] ResNet-50-FPN 37.9 60.6 40.9 23.0 46.8 52.7 6.1
R-ARE-Net [37] DarkNet-19 38.0 60.8 40.7 25.2 45.9 52.7 -
RB Mask R-CNN [6] ViT-Tiny 38.5 61.5 41.4 24.0 47.3 53.3 3.8
ISOP [9] ResNet-101-FPN 39.4 64.2 41.7 24.9 48.6 55.4 16.6
Ye et al. [36] Swin-Tiny 40.2 65.1 44.3 25.3 49.9 57.3 2.9
ISOP [9] ViT-Tiny 40.3 65.2 43.9 25.6 49.4 56.8 9.1
Finetuned MobileSAM [17] ViT-Tiny 33.4 58.8 32.9 19.1 42.5 47.8 21.5
Finetuned EfficientSAM [16] ViT-Tiny 34.8 60.0 34.0 20.3 43.7 48.6 21.5
Finetuned SAM [10] ViT-H 37.2 62.3 36.7 22.9 45.9 51.1 2.8
RSPrompter-anchor [11] ViT-H 40.9 64.0 44.2 24.3 49.7 57.0 2.7
RSPrompter-query [11] ViT-H 41.6 64.4 44.6 24.8 50.2 57.6 2.7
OBSeg* ViT-H 45.1 69.4 48.3 28.8 55.1 60.7 2.7
[…]
OBSeg ViT-Tiny 43.7 65.2 46.1 25.7 51.4 59.6 20.4
@@ IV Experiments
To verify the effectiveness of OBSeg, we conducted experiments on multiple public datasets. Since some datasets do not provide the GT parameters of OBBs, the smallest OBBs of instance masks are used as regression targets.
@@ IV-A Dataset
IV-A1 iSAID
iSAID [38] is a large-scale and densely annotated instance segmentation dataset in aerial images. It contains 2806 high-resolution aerial images and 655,451 object instances across 15 categories. The categories include airplanes, ships, storage tanks, baseball fields, tennis courts, basketball courts, ground track and field, ports, bridges, large vehicles, small vehicles, helicopters, roundabouts, swimming pools, and football fields. The training set, validation set and test set contain 1411, 458 and 937 images, respectively. All images are cropped into $1024\times 1024$ with a stride of 500.
IV-A2 NWPU VHR-10
NWPU VHR-10 instance segmentation dataset [39] is a widely-used dataset in geospatial tasks, which contains 800 high-resolution color and panchromatic infrared images. It has 650 very-high-resolution (VHR) optical images with objects and 150 VHR optical images with pure background. We randomly divide the dataset into the training set and the test set in a ratio of $7\colon 3$.
IV-A3 PSeg-SSDD
PSeg-SSDD dataset [40] supports ship instance segmentation and ship detection tasks. It contains 1160 SAR images with a resolution ranging from 1 to 15m, and there are various polarizations, resolutions, and scenes. We randomly divide this dataset into a training set of 928 images and a test set of 232 images.
Figure 6: Some comparison results on the iSAID dataset. Segmentation masks of different colors represent different instances.
@@ IV-B Implementation Details
In this work, OBSeg chooses SAM as the segmentation foundation model. For OBB detection, we directly employ Oriented R-CNN [12], which is a two-stage OBB detection model with competitive detection performance. The training settings for Oriented R-CNN with ResNet-34 [41] are the same as the original method. In the OBB prompt encoder, the length of the sampled Fourier basis frequencies is 128. For knowledge distillation, the teacher model (called OBSeg*) is based on SAM with the powerful ViT-H (vision transformer [42]) image encoder and the proposed OBB prompt encoder. The student model (i.e., OBSeg) directly uses the pretrained lightweight ViT-Tiny image encoder in MobileSAM and the proposed OBB prompt encoder. Their mask decoders and OBB detection modules are the same. During training, the weights of their image encoders are frozen, and the other parts are fine-tuned. Fine-tuning and distillation are performed for 5 and 8 epochs, respectively. The teacher model has the same training loss as SAM and has been trained before distillation. The trade-off factor $\lambda$ for the distillation loss is 0.1. More implementation details can be found in our open-source code. All experiments are conducted on an RTX 3090 GPU. For evaluation, we use the standard evaluation metrics: AP (i.e., AP50-95), AP50, AP75, APS, APM and APL, where “AP” means average precision over IoU (Intersection over Union) threshold.
TABLE II: Comparison results on the NWPU VHR-10 test set
Method Backbone AP AP50 AP75 APS APM APL
YOLACT [43] ResNet-50-FPN 43.3 77.5 44.0 23.1 40.5 54.3
Mask Scoring R-CNN [32] ResNet-50-FPN 58.4 92.1 63.2 46.7 56.9 63.5
PointRend [31] ResNet-50-FPN 61.1 90.1 64.7 52.8 59.9 61.0
PointRend [31] ViT-Tiny 61.9 91.2 65.8 53.4 61.7 61.6
Hybrid Task Cascade [35] ResNet-50-FPN 62.8 92.9 69.1 49.8 63.5 65.8
ARE-Net [37] ResNet-101-FPN 64.8 93.2 71.5 53.9 65.3 72.9
Shi and Zhang [44] ResNet-FPN 65.2 94.9 72.1 49.4 65.7 71.2
FB-ISNet [45] DLA-BiFPN 67.0 94.5 72.0 - - -
DCTC [33] DLASeg 67.7 91.8 75.5 56.4 66.4 69.9
Ye et al. [36] Swin-Tiny 68.2 - - 50.9 68.4 65.7
Ye et al. [36] ViT-Tiny 68.2 - - 51.4 68.0 65.8
Finetuned MobileSAM [17] ViT-Tiny 60.8 91.0 63.1 51.4 61.5 60.7
Finetuned EfficientSAM [16] ViT-Tiny 62.6 92.5 65.9 51.8 63.0 63.1
Finetuned SAM [10] ViT-H 65.0 95.0 70.4 53.6 65.2 73.0
RSPrompter-anchor [11] ViT-H 66.1 92.7 70.6 - - -
RSPrompter-query [11] ViT-H 67.5 91.7 74.8 - - -
OBSeg* ViT-H 71.1 96.7 77.4 59.0 70.5 74.6
@@
OBSeg ViT-Tiny 69.5 95.1 75.6 57.7 69.2 73.3
TABLE III: Comparison results on the PSeg-SSDD test set
Method Backbone AP AP50 AP75 APS APM APL
YOLACT [43] ResNet-50-FPN 57.8 91.4 70.9 58.4 56.5 58.6
SA R-CNN [46] ResNet50-FPN 59.4 90.4 77.6 63.3 62.5 47.7
EMIN [47] - 61.7 94.3 76.8 62.1 61.3 40.1
FL-CSE-ROIE [48] ResNet-101-FPN 62.6 93.7 78.3 63.3 61.2 75.0
MAI-SE-Net [49] ResNet-101-FPN 63.0 94.4 77.6 63.3 62.5 47.7
LFG-Net [50] ResNeXt-64 64.2 95.0 81.1 63.1 68.2 43.1
Mask Scoring R-CNN [32] ResNet-50-FPN 65.4 94.1 81.9 66.4 62.0 40.5
PointRend [31] ResNet-50-FPN 65.6 94.5 82.3 67.0 62.2 16.8
Hybrid Task Cascade [35] ResNet-50-FPN 66.1 95.4 83.2 66.2 68.9 29.4
PointRend [31] ViT-Tiny 66.2 95.2 83.3 67.7 64.0 20.4
DCTC [33] DLASeg 67.4 93.9 83.8 64.4 76.6 80.2
Shi and Zhang [44] ResNet-FPN 67.8 97.5 85.0 67.7 69.1 58.9
Ye et al. [36] Swin-Tiny 68.4 - - 69.8 63.7 41.0
Ye et al. [36] ViT-Tiny 68.5 - - 69.4 64.6 42.5
Finetuned MobileSAM [17] ViT-Tiny 62.9 93.6 77.1 64.0 64.8 49.9
Finetuned EfficientSAM [16] ViT-Tiny 63.5 94.5 79.6 64.3 65.2 54.3
Finetuned SAM [10] ViT-H 66.5 95.0 83.3 64.4 66.7 58.2
RSPrompter-anchor [11] ViT-H 66.8 94.7 84.0 - - -
RSPrompter-query [11] ViT-H 67.3 95.6 84.3 - - -
OBSeg* ViT-H 71.8 97.0 87.5 71.2 75.3 72.4
@@
OBSeg ViT-Tiny 69.8 96.1 85.4 69.4 73.1 66.6
IV-C Comparisons with the State-of-the-Art
IV-C1 Comparison Experiments on the iSAID Dataset
To evaluate the effectiveness of OBSeg, we compare OBSeg with current instance segmentation methods using OBBs on the iSAID dataset. We also include some instance segmentation methods using HBBs (e.g., SAM after finetuned on this dataset) and other SAM-based methods using HBBs (e.g., RSPrompter) for comparison. For fairness, SAM and the teacher model OBSeg* use ViT-H as the image encoder, and OBSeg uses ViT-Tiny as the image encoder. We also replace the backbones of some non-BSM-based methods with pretrained ViT-Tiny. SAM, MobileSAM, and EfficientSAM apply powerful YOLOv7 [51] as the HBB detector. $\dagger$” indicates that Gaussian smoothing is not applied during distillation, similar to its use in current lightweight BSMs. The inference is conducted on the RTX 3090 GPU, and the inference speed of some models is based on the results we reproduced.
As presented in Table I, OBSeg* and OBSeg significantly outperform other methods, achieving 45.1% AP and 43.7% AP, respectively. Compared with instance segmentation methods using HBBs, methods using OBBs perform better since OBBs are more suitable for segmenting densely packed objects. Significant accuracy improvements over fine-tuned SAM series and other SAM-based methods using HBBs also indicate the benefit of using OBBs. In addition, OBSeg is less dependent on OBB detection performance and outperforms existing “segmentation within bounding box” methods by a large margin. Fig. 6 shows some instance segmentation visualization results of OBSeg and current methods. For dense objects with oblique orientations, the segmentation results of methods using HBBs incorrectly include other instances, reducing the instance segmentation performance. Furthermore, when predicted OBBs significantly deviate from GT OBBs, the segmentation performance of existing “segmentation within bounding box” methods is substantially affected. Based on the OBB prompt mechanism, OBSeg still accurately segments instances when the OBB predictions are inaccurate.
Compared with other BSMs, OBSeg is very lightweight. Even though SAM has a powerful ViT-H image encoder, its performance is still inferior to OBSeg. Among BSMs of the same lightweight level, OBSeg achieves higher accuracy and a competitive inference speed of 20.4 FPS, which is also better than most methods not based on BSMs. This demonstrates that OBSeg is more efficient than current instance segmentation methods. Moreover, compared with OBSeg*, the notable inference speed improvement and minor accuracy loss of OBSeg also demonstrate the effectiveness of the proposed knowledge distillation method. Gaussian smoothing further enhances the distillation performance and generalization capability. More instance segmentation results of OBSeg on the iSAID dataset are displayed in Fig. 5.
IV-C2 Comparison Experiments on the NWPU VHR-10 Dataset
We further verify the instance segmentation performance of OBSeg on the NWPU VHR-10 dataset. For fairness, we compare instance segmentation methods using HBBs, methods using OBBs, and BSM-based methods. As in the setting on the iSAID dataset, the HBB detectors in the BSM series all use the powerful YOLOv7. We also replace the backbones of some non-BSM-based methods with ViT-Tiny.
[…]
IV-C3 Comparison Experiments on the PSeg-SSDD Dataset
Similar to the experimental settings of the iSAID and NWPU VHR-10 datasets, we compare methods that use HBBs, OBBs, and those based on BSMs for instance segmentation.
As presented in Table III, similar to the comparison experiment results on the iSAID and NWPU VHR-10 datasets, OBSeg* and OBSeg show their superior performance. Before distillation, the teacher model OBSeg* achieves an accuracy of 71.8% AP. After distillation, OBSeg greatly retains the model’s performance, achieving an instance segmentation accuracy of 69.8% AP, which is superior to current methods.
IV-D Ablation Studies
@@
To evaluate the effectiveness of the components of OBSeg, a series of ablation experiments are conducted on the iSAID dataset. We use the most commonly used AP50 as the metric.
IV-D1 Oriented Bounding Box Prompt Encoder
@@
The OBB prompt encoder introduces positional encoding and specific embeddings to encode OBBs. In Table IV, “OCE”, “TCE” and “BCE” represent the specific orientation correction embeddings, top-left corner embeddings and bottom-right corner embeddings, respectively. Compared with GPE, AGPE encodes location information better by adaptively learning parameters of Gaussian distributions. Moreover, specific embeddings further improve instance segmentation performance by 0.97% AP50. Orientation correction embeddings alleviate the periodicity problem of orientation.
IV-D2 Knowledge Distillation
We compare the performance of OBSeg learning from the teacher model and GT labels, and test the impact of different label smoothing methods. “T” represents learning from the teacher model. The smoothing factor in ULS is 0.1. For distillation on the mask decoder, Gaussian smoothing (GS) has the same settings as SVLS, with a Gaussian kernel size of $3\times 3$ and a standard deviation of 1.0. For distillation on the OBB prompt encoder, GS has a Gaussian kernel length of 3 and a standard deviation of 0.3. The weight factor of the teacher model and GT labels is 0.1. “GS-” represents distillation only for the mask decoder.
The experimental results in Table V demonstrate that learning from the teacher model has higher performance than learning from GT labels (learning only from GT labels means no knowledge distillation is performed), which is mainly thanks to the powerful data-fitting capabilities of BSMs. Furthermore, adding distillation for the OBB prompt encoder effectively enhances performance. Moreover, compared with distillation methods that are not based on Gaussian smoothing in existing lightweight distilled BSMs (e.g., EfficientSAM and MobileSAM), Gaussian smoothing in OBSeg further improves distillation performance and is superior to ULS. In addition, we compare the impact of different parameter settings in Gaussian smoothing on instance segmentation performance in Table VI. When $k=5$, $\sigma=0.3$, and $\delta=1.0$, the highest instance segmentation accuracy of 65.24% AP50 is achieved.
TABLE IV: Performance of each component of the OBB prompt encoder
@@ IV-D2 Knowledge Distillation
64.00
64.27
64.86
65.24
TABLE V: Comparison of different learning targets and label smoothing methods in knowledge distillation
@@
64.23 64.44 64.73 64.79 64.87 64.91 65.16
TABLE VI: Impact of parameter settings in Gaussian smoothing
@@
$k=3$ 65.16 65.16 65.12 65.09
$k=5$ 65.14 65.24 65.10 65.13
$k=7$ 65.11 65.17 65.07 65.06
TABLE VII: Performance comparison of OBB and HBB detectors
YOLOv7 [51] Oriented Oriented S2ANet [52]
R-CNN [12] Reppoints [53]
58.81 65.24 65.37 65.09
IV-D3 Oriented Bounding Box Detection